# A Statistical Framework for Understanding Causal Effects that Vary by Treatment Initiation Time in EHR-based Studies

### v1

@@ 3.2 Calendar Time-Specific Average Treatment Effects
Notice that $\chi_{m}(P)$ is simply the calendar time-specific ATE among patients eligible for trial $m$. Implicit in the definition of $\chi_{m}(P)$ is that comparisons are being made between the specific versions of treatment (i.e., surgical technique, standard of care) in place at time $m$. Note, $\chi_{m}(P)$ is an intention-to-treat (ITT) effect (hernan2016target) in the sense that it captures the effect of initiating treatment at time $m$ $(a_{m}=1)$ versus not initiating $(a_{m}=0)$ without regard to what happens at future time points. In particular, the definition of $\chi_{m}(P)$ is not concerned that treatment definitions may evolve over time (e.g., $a_{j}=0$ and $a_{m}=0$ may mean slightly different things for $j>m$), nor for the possibility of treatment switching.
@@ 4 Estimation of Calendar Time-Specific Average Treatment Effects
In our data application, we take $\bm{L}_{m}$ to be the most recent values of covariates outlined in Figure 1, although we note that analysts can also include in this vector appropriate summaries of covariate and/or treatment history prior to baseline of trial $m$, if warranted in their application.
@@ 4.1 Motivating A Projection Parameter Approach
At the outset, estimation of each calendar time-specific ATE $\chi_{m}(P)$ may seem straightforward. Let $\mu_{m}(a_{m},\bm{L}_{m})=\mathbb{E}[Y_{m}\penalty 10000\ |\penalty 10000\ A_{m}=a_{m},\bm{L}_{m},E_{m}=1]$ and $\pi_{m}(L_{m})=P(A_{m}=1\penalty 10000\ |\penalty 10000\ \bm{L}_{m},E_{m}=1)$ denote the trial $m$ specific outcome regression and propensity score models, respectively. Under Assumptions 1-3, $\chi_{m}(P)$ is identified by the $g$-formula, $\mathbb{E}[\mu_{m}(1,\bm{L}_{m})-\mu_{m}(0,\bm{L}_{m})\penalty 10000\ |\penalty 10000\ E_{m}=1]$ or via inverse probability weighting, $\mathbb{E}\big[\big(\frac{A_{m}}{\pi_{m}(\bm{L}_{m})}-\frac{1-A_{m}}{1-\pi_{m}(\bm{L}_{m})}\big)Y_{m}\penalty 10000\ |\penalty 10000\ E_{m}=1\big]$ (hernan2024). One might then proceed to estimate $\chi_{m}(P)$ on the basis of its nonparametric influence function (tsiatis_semiparametric_2006, kennedy2023semiparametric),
@@
Unfortunately, estimating $\mu_{m}$ or $\pi_{m}$ by fitting a separate model for each trial would be difficult in practice, especially in settings where a relatively small number of patients initiate a particular treatment during trial $m$. For example, 24 of 84 target trials in our study have fewer than 100 patients undergoing bariatric surgery at the corresponding baseline month, and 12 of 84 of have fewer than 25 surgical patients. This problem is further compounded for comparisons involving the individual surgical procedures, RYGB and SG (Figure S2).
The predominant analytic strategy in sequential target trial emulations for time-to-event outcomes involves fitting a single parametric outcome and/or propensity score model pooling information across trials (scola2023implementation) (e.g., pooled logistic regression (dagostino1990)), perhaps with a spline term to flexibly adjust for calendar time as a confounder (dickerman2019, madenci2024). Given that trends in $\chi_{m}(P)$ over calendar time may reflect a complex interplay between changing treatment efficacy and shifts in effect modifier distributions, applying a pooled parametric model in this setting would require correctly specifying a substantial number of flexible interactions between treatment, confounders, and calendar time. Such a task may be unrealistic in many instances—arguably including our data application given the heterogeneity reported in Figure 1—and might instead motivate the use of pooled models based on flexible machine learning methods.
Ultimately, while we are interested in trial-specific ATEs, we are most interested in the underlying trend in $\chi_{m}(P)$ across calendar time. To formalize estimation of this trend across calendar time, we follow the strategy of kennedy2019 and others (kennedy2023density, neugebauer2007nonparametric, rosenblum2010targeted), and formulate our estimand as a projection of all $M$ trial-specific effects onto a candidate parametric model. To operationalize this approach, we first define a MSM for mean trial-specific counterfactual outcomes.
@@
When the MSM in Equation (1) is saturated, $\psi(m)=\chi_{m}(P)$. Instead, we adopt a parametric model for $\psi(m)$, denoted $\psi(m;\bm{\beta})$, and take our estimand to be the projection of $\chi_{m}(P)$ onto $\psi(m;\bm{\beta})$. Formally, $\bm{\beta}(P)$ is the projection parameter defined by
$\bm{\beta}(P)=\mathop{\mathrm{argmin}}_{\bm{\beta}\in\mathbb{R}^{k}}\sum_{m=1}^{M}h(m)\Bigr(\chi_{m}(P)-\psi(m;\bm{\beta})\Bigr)^{2},$
where $h(m)$ is an optional trial-specific weight. Henceforth, we proceed with weights of the form $h(m)=w(m)P(E_{m}=1)$ to weight trial contributions according to their effective sample size, where $w(m)$ allows for optional weighting by other additional factors.
This approach offers several advantages, which we elaborate on in subsequent sections. As will become clear in Section 4.2, estimation of $\chi_{m}(P)$ is a necessary step towards estimating $\psi(m;\bm{\beta})$, a task for which we advocate the use of pooled nuisance models based on flexible machine learning methods. Nevertheless, our projection parameter approach allows us to identify meaningful underlying trends in $\chi_{m}(P)$ even if some trial-specific effects are estimated imprecisely (e.g., due to small sample sizes). Although pooled nonparametric models for component nuisance functions will be an essential step in estimating $\chi_{m}(P)$, they provide limited insight into how information is shared across calendar time. Introducing the additional projection step makes this information sharing explicit and clarifies how effect estimates from individual trials are combined to learn trends over calendar time, thereby improving transparency relative to relying on pooled nonparametric models alone. Finally, this approach naturally supports a model selection procedure (described in Section 4.4) for comparing candidate MSMs to identify the functional form that best describes how treatment effects vary over calendar time, including the possibility of a single, constant effect.
@@ 4.2 Efficient Estimation of Projection Parameter
We now introduce preliminary efficiency theory for the projection parameters $\bm{\beta}(P)$, which will motivate our proposed estimator. Since influence functions are by definition mean zero, it is helpful to work with uncentered versions, $\dot{\chi}_{m}(O;P)\coloneqq\frac{\mathds{1}(E_{m}=1)}{P(E_{m}=1)}\dot{\chi}^{\dagger}_{m}(O;P)$, which satisfy $\mathbb{E}_{P}[\dot{\chi}_{m}(O;P)]=\chi_{m}(P)$, where
@@
These quantities play an important role in the influence function of $\bm{\beta}(P)$, as summarized in the following theorem.
@@
Suppose that the candidate MSM, $\psi(m;\bm{\beta})$ is twice differentiable in $\bm{\beta}$. Then, under Assumptions 1-3, $\bm{\beta}(P)$ is identified as the solution to the estimating equation,
@@
where $\widehat{\chi}_{m}$ is the influence function-based estimator of $\chi_{m}(P)$ mentioned in Section 4.1, and $\widehat{P}$ indicates that any necessary nuisance functions are replaced by estimated quantities. Equations (2) and (3) make clear, however, that to even estimate $\bm{\beta}(P)$, estimates of $\chi_{m}(P)$, and thus estimates of nuisance functions $\mu_{m}$ and $\pi_{m}$, are required. Also of note in Equation (3), trials are weighted proportionally to the sample size of their eligible populations, potentially mitigating the influence of noisier estimates of $\chi_{m}(P)$ in component trials with smaller sample sizes.
@@
In order to estimate $\chi_{m}(P)$, we first create a pooled dataset $\mathcal{D}$ of patient information across all trials in which they are eligible, that is,
@@
Routines for estimation of influence function-based estimators like $\widehat{\chi}_{m}$ or $\widehat{\bm{\beta}}$ and typically involve sample splitting or cross-fitting, particularly when using nonparametric and/or machine learning techniques to estimate component nuisance functions (chernozhukov2018). We describe the complete procedure for estimating $\widehat{\chi}_{m}$ and $\widehat{\bm{\beta}}$ below.
@@
    For a given candidate marginal structural model, $\psi(m;\bm{\beta})$, plug estimates $\{\widehat{\chi}_{m}\}_{m=1}^{M}$ into Equation (3) and solve for $\widehat{\bm{\beta}}$ (e.g., via nleqslv (R-nleqslv)).
@@ 4.3 Theoretical Properties of $\widehat{\bm{\beta}}$ and $\widehat{\chi}_{m}$
Pooled models can help ensure that the required rate conditions are met in component trials with smaller effective sample sizes. As discussed in Section 4.2, however, pooled parametric models for $\pi_{m}$ and $\mu_{m}$ are likely to be inconsistent in the presence of significant covariate shift, motivating the use of more flexible pooled models. Crucially, when the required rate conditions (e.g., $o_{P}(n^{-1/4})$ for each nuisance function) are satisfied, asymptotically valid inference for the trend in treatment effects across calendar time remains attainable, even when using such flexible pooled models.
In particular, asymptotically valid pointwise Wald-style $(1-\alpha)$-level confidence intervals are given by
@@
4.4 Model Selection Among Candidate Marginal Structural Models
Suppose one considers a range of candidate MSMs including a constant model (i.e., $\psi(m;\beta)=\beta$), a linear model (i.e., $\psi(m;\bm{\beta})=\beta_{0}+\beta_{1}m$), and more flexible MSMs including higher-order polynomials (in $m$) and splines. A model selection procedure to select the MSM which best describes the underlying trend $\chi_{m}(P)$ is important to best characterize how treatment effects—and thus possibly treatment efficacy—vary over calendar time. In particular, if a constant model were selected, it would suggest that any underlying variation in treatment effects is not detectable given the study sample size, and thus may not be clinically meaningful. On the other hand, if a flexible MSM were selected, it would provide evidence of non-constant trends in treatment effects over calendar time.
To implement such a model selection procedure, we adapt the ideas of kennedy2019 and vanderLaanDudoit2003cv. Let $\{\widehat{\psi}_{1},...,\widehat{\psi}_{K}\}$ denote a set of $K$ candidate MSM projection models, with $\widehat{\psi}_{k}\equiv\psi_{k}(m;\widehat{\bm{\beta}}_{k})$. A natural quantity by which to compare candidate MSMs is mean-squared error (relative to the true $\chi_{m}(P)$). To simplify comparisons, we drop terms that don’t depend on the candidate MSM and proceed with the pseudorisk,
@@ 4.4 Model Selection Among Candidate Marginal Structural Models
This is an intuitive loss function to work with because it reflects an estimating equation-based estimator of the pseudorisk, based on its influence function, treating $\widehat{\psi}_{k}$ as fixed (kennedy2019).
@@
In practice, differences between candidate projection models can be small (kennedy2023semiparametric), particularly if models classes are nested within one another. In such cases, we might have a preference for simpler models, such that more flexible MSMs are selected only if there are clinically meaningful trends in $\chi_{m}(P)$ across calendar time. To formalize this, let $\widehat{\psi}^{*}=\mathop{\mathrm{argmin}}_{\widehat{\psi}_{k}\in\{\widehat{\psi}_{1},\dots,\widehat{\psi}_{K}\}}\widehat{L}(\widehat{\psi}_{k})$ denote the MSM among our set of candidates which minimizes the loss function $\widehat{L}(\widehat{\psi}_{k})$. Let
@@
One can interpret $\epsilon$ as a pooled standard deviation across our best estimate of the calendar time trend, weighting proportional to the size of the eligible population at each trial. Finally, one can choose the simplest model such that $|\widehat{L}(\widehat{\psi}_{k})-\widehat{L}(\widehat{\psi}^{*})|\leq c\epsilon$, for some user-selected constant $c$. There is considerable debate in the literature about how large an effect size needs to be on a standard deviation scale to be clinically meaningful (tsujimoto2022mic), though the notion of clinically meaningful is very application dependent. In simulations, we find that $c=1/4$ works well in our setup, but present result for a range of other values. Finally, we note that notion of “simpler model” can be operationalized through any model complexity criterion so long as models can be ordered by this criterion. One reasonable such choice—which we adopt in our application—is ordering candidate MSMs by the dimension of their indexing projection parameter (i.e., the dimension of $\bm{\beta}$).

### v2

@@ 3.2 Calendar Time-Specific Average Treatment Effects
Notice that $\chi_{m}(P)$ is simply the calendar time-specific ATE among patients eligible for trial $m$. Implicit in the definition of $\chi_{m}(P)$ is that comparisons are being made between the specific versions of treatment (i.e., surgical technique, standard of care) in place at time $m$. We are not interested in defining nor estimating effects of dynamic treatment regimes and note that consistent with the definition of $A_{m}$ as a point exposure, $\chi_{m}(P)$ captures the effect of initiating treatment at time $m$ $(a_{m}=1)$ versus not initiating $(a_{m}=0)$ without regard to what happens at future time points. In particular, the definition of $\chi_{m}(P)$ is not concerned that treatment definitions may evolve over time (e.g., $a_{j}=0$ and $a_{m}=0$ may mean slightly different things for $j>m$), nor for the possibility of treatment switching; see Section 8 for a discussion of how our framing relates to the consistency assumption and “multiple versions of treatment”. Due to the rarity of initiating surgery relative to standard of care and since bariatric surgery is effectively an irreversible procedure (at least in so far as it relates to a patient adhering to surgery or standard of care), non-adherence is extremely uncommon in our study.
Just as a constant hazard ratio represents a weighted average of time-varying hazard ratios (across time since treatment initiation) (Hernán, 2010), a sequential target trial emulation estimating a constant ATE across calendar time implicitly targets some weighted average of trial-specific effects $\chi_{m}(P)$. However, the precise definition and causal interpretation of such contrast is rarely formalized or made transparent.
@@ 4 Estimation of Calendar Time-Specific Average Treatment Effects
In order to answer the clinical questions raised in Section 2, one might initially target the nonparametric curve composed of the collection $\bigr\{\chi_{m}(P)\}_{m=1}^{M}$. As we elaborate upon, however, a small number of surgical patients in any particular target trial presents a challenge towards efficient estimation of any individual $\chi_{m}(P)$. MSMs have proven useful for analysis of continuous exposures by parameterizing a dose-response curve to make causal inference tractable Robins et al. (2000), Naimi et al. (2014), Hernán and Robins (2024). In that same vein, pragmatic adoption of MSMs in the present context provides structure on top of $\bigr\{\chi_{m}(P)\}_{m=1}^{M}$, allowing information to be borrowed across target trials to describe trends over calendar time.
4.1 Assumptions
@@
Following closely to Arterburn et al. (Arterburn et al., 2020), we take $\bm{L}_{m}$ to be the most recent values of covariates outlined in Figure 1, although we note that analysts can also include in this vector appropriate summaries of covariate and/or treatment history prior to baseline of trial $m$, if warranted in their application. A careful discussion of Assumption 1, particularly in the context of multiple versions of treatment and changing treatment efficacy, is deferred until Section 8.
@@ 4.2 Motivating A Projection Parameter Approach
Let $\mu_{m}(a_{m},\bm{L}_{m})=\mathbb{E}[Y_{m}\penalty 10000\ |\penalty 10000\ A_{m}=a_{m},\bm{L}_{m},E_{m}=1]$ and $\pi_{m}(L_{m})=P(A_{m}=1\penalty 10000\ |\penalty 10000\ \bm{L}_{m},E_{m}=1)$ denote the trial $m$-specific outcome regression and propensity score models, respectively. Under Assumptions 1-3, $\chi_{m}(P)$ is identified by the $g$-formula, $\mathbb{E}[\mu_{m}(1,\bm{L}_{m})-\mu_{m}(0,\bm{L}_{m})\penalty 10000\ |\penalty 10000\ E_{m}=1]$ or via inverse probability weighting, $\mathbb{E}\big[\big(\frac{A_{m}}{\pi_{m}(\bm{L}_{m})}-\frac{1-A_{m}}{1-\pi_{m}(\bm{L}_{m})}\big)Y_{m}\penalty 10000\ |\penalty 10000\ E_{m}=1\big]$ (Hernán and Robins, 2024). One might then proceed to estimate $\chi_{m}(P)$ on the basis of its nonparametric influence function (Tsiatis, 2006, Kennedy, 2023),
@@
Unfortunately, estimating $\mu_{m}$ or $\pi_{m}$ by fitting a separate model for each trial would be difficult in practice, especially in settings where a relatively small number of patients initiate a particular treatment during trial $m$. For example, 24 of 84 target trials in our study have fewer than 100 patients undergoing bariatric surgery at the corresponding baseline month, and 12 of 84 have fewer than 25 surgical patients. In fact, the lack of power from small sample sizes inherent to emulating a target trial from a single enrollment window was one of the primary motivations behind development of the sequential target trial design (Hernán et al., 2008).
The predominant analytic strategy in sequential target trial emulations for time-to-event outcomes targets (discrete time) hazard ratios by fitting a single parametric outcome and/or propensity score model pooling information across trials (Scola et al., 2023) (e.g., pooled logistic regression (D’Agostino et al., 1990)), perhaps with spline terms to enable non-proportional hazards and flexibly adjust for calendar time as a confounder (Dickerman et al., 2019, Madenci et al., 2024). Given that trends in $\chi_{m}(P)$ over calendar time may reflect a complex interplay between changing treatment efficacy and shifts in effect modifier distributions, applying a pooled parametric model in this setting would require correctly specifying a substantial number of flexible interactions between treatment, confounders, and calendar time. Such a task may be unrealistic in many instances—arguably including our study of bariatric surgery given the heterogeneity reported in Figure 1—and might instead motivate the use of pooled models based on flexible machine learning methods.
Ultimately, while we are interested in trial-specific ATEs, we are most interested in characterization of the underlying trend in $\chi_{m}(P)$ across calendar time. To formalize estimation of this trend across calendar time, we follow the strategy of Kennedy et al. Kennedy et al. (2019) and others (Kennedy et al., 2023, Neugebauer and Van der Laan, 2007, Rosenblum and Van der Laan, 2010), and formulate our estimand as a projection of all $M$ trial-specific effects onto a candidate parametric marginal structural model. Specifically, following the pragmatic use of MSMs for analysis of continuous exposures (Kennedy et al., 2019, Naimi et al., 2014, Hernán and Robins, 2024, Robins et al., 2000), we define a MSM for mean trial-specific counterfactual outcomes.
@@
When the MSM in Equation (1) is saturated, $\psi(m)=\chi_{m}(P)$, that is, we are targeting the true nonparametric effect curve across calendar time. Because such contrast is not easily estimable due to challenges described above, we adopt a working parametric model for $\psi(m)$, denoted $\psi(m;\bm{\beta})$, and take our estimand to be the projection of $\chi_{m}(P)$ onto $\psi(m;\bm{\beta})$. Despite being a working model, $\psi(m;\bm{\beta})$ remains a useful summary of how treatment efficacy has evolved over calendar time. Moreover, when study data are proximal to the present, the more recent end of this curve, $\psi(M;\bm{\beta})$, offers direct insight into treatment efficacy in current clinical practice.
Formally, $\bm{\beta}(P)$ is the projection parameter defined by
$\bm{\beta}(P)=\mathop{\mathrm{argmin}}_{\bm{\beta}\in\mathbb{R}^{k}}\sum_{m=1}^{M}w(m)P(E_{m}=1)\Bigr(\chi_{m}(P)-\psi(m;\bm{\beta})\Bigr)^{2},$
where $P(E_{m}=1)$ weights trial contributions according to their effective sample size and $w(m)$ allows for optional weighting by other additional factors. We emphasize that $\psi(m;\bm{\beta})$ is not believed to be the true effect curve but rather a tractable summary of trends across calendar time. Because the form of $\psi(m;\bm{\beta})$ that best approximates the nonparametric curve $\bigr\{\chi_{m}(P)\}_{m=1}^{M}$ is unknown, we outline a model selection procedure (described in Section 4.5) for comparing candidate MSMs to identify the functional form that best describes how treatment effects vary over calendar time, including the possibility of a single, constant effect.
@@ 4.3 Efficient Estimation of Projection Parameter
We now introduce preliminary efficiency theory for the projection parameters $\bm{\beta}(P)$, which will motivate our proposed estimator. Since influence functions are by definition mean zero, it is helpful when constructing estimators to work with uncentered versions, $\dot{\chi}_{m}(O;P)\coloneqq\frac{\mathds{1}(E_{m}=1)}{P(E_{m}=1)}\dot{\chi}^{\dagger}_{m}(O;P)$, which satisfy $\mathbb{E}_{P}[\dot{\chi}_{m}(O;P)]=\chi_{m}(P)$, where
@@
Suppose that the candidate MSM, $\psi(m;\bm{\beta})$ is twice differentiable in $\bm{\beta}$. Under Assumptions 1-3, $\bm{\beta}(P)$ is identified as the solution to the estimating equation,
@@
where $\widehat{\chi}_{m}$ is the influence function-based estimator of $\chi_{m}(P)$ mentioned in Section 4.2, and $\widehat{P}$ indicates that any necessary nuisance functions are replaced by estimated quantities. Equations (2) and (3) make clear, however, that to even estimate $\bm{\beta}(P)$, estimates of $\chi_{m}(P)$, and thus estimates of nuisance functions $\mu_{m}$ and $\pi_{m}$, are required.
@@
To estimate $\chi_{m}(P)$, we first create a pooled dataset $\mathcal{D}$ of patient information across all trials in which they are eligible, that is,
@@
Routines for estimation of influence function-based estimators like $\widehat{\chi}_{m}$ or $\widehat{\bm{\beta}}$ typically involve sample splitting or cross-fitting, particularly when using nonparametric and/or machine learning techniques to estimate component nuisance functions (Chernozhukov et al., 2018). We describe the complete procedure for estimating $\widehat{\chi}_{m}$ and $\widehat{\bm{\beta}}$ in Algorithm 1 below.
@@
    For a given candidate marginal structural model, $\psi(m;\bm{\beta})$, plug estimates $\{\widehat{\chi}_{m}\}_{m=1}^{M}$ into Equation (3) and solve for $\widehat{\bm{\beta}}$ (e.g., via nleqslv (Hasselman, 2023)).
@@ 4.4 Theoretical Properties of $\widehat{\bm{\beta}}$ and $\widehat{\chi}_{m}$
Practically, pooled models can help ensure that the required rate conditions are met in component trials with smaller effective sample sizes. As discussed in Section 4.3, however, pooled parametric models for $\pi_{m}$ and $\mu_{m}$ are likely to be inconsistent in the presence of significant covariate shift, motivating the use of more flexible pooled models. Crucially, when the required rate conditions (e.g., $o_{P}(n^{-1/4})$ for each nuisance function) are satisfied, asymptotically valid inference for the trend in treatment effects across calendar time remains attainable, even when using such flexible pooled models. In particular, asymptotically valid pointwise Wald-style $(1-\alpha)$-level confidence intervals are given by
@@
4.5 Selection Among Candidate Marginal Structural Models
Unless an analyst a priori desires to summarize treatment efficacy through a single constant effect across calendar time, it is often unclear how to specify the functional form for $\psi(m;\bm{\beta})$ which appropriately summarizes the nonparametric curve $\bigr\{\chi_{m}(P)\}_{m=1}^{M}$, and in practice several candidate models may be reasonable. Forcing a single specification risks either under- or over-smoothing the underlying trend in $\chi_{m}(P)$, obscuring clinically meaningful variation or introducing spurious structure. It is therefore preferable to consider a range of candidate MSMs, including a constant model (i.e., $\psi(m;\beta)=\beta$), a linear model (i.e., $\psi(m;\bm{\beta})=\beta_{0}+\beta_{1}m$), and more flexible MSMs including higher-order polynomials (in $m$) and splines, and select among them using a data-adaptive procedure that lets the data inform the appropriate level of complexity.
To implement such an MSM selection procedure, we adapt the ideas of Kennedy et al.Kennedy et al. (2019) and van der Laan and Dudoit van der Laan and Dudoit (2003). Let $\{\widehat{\psi}_{1},...,\widehat{\psi}_{K}\}$ denote a set of $K$ candidate MSM projection models, with $\widehat{\psi}_{k}\equiv\psi_{k}(m;\widehat{\bm{\beta}}_{k})$. A natural quantity by which to compare candidate MSMs is mean-squared error (relative to the true $\chi_{m}(P)$). To simplify comparisons, we drop terms that do not depend on the candidate MSM and proceed with the pseudorisk,
@@ 4.5 Selection Among Candidate Marginal Structural Models
This is an intuitive loss function to work with because it reflects an estimating equation-based estimator of the pseudorisk, based on its influence function, treating $\widehat{\psi}_{k}$ as fixed (Kennedy et al., 2019).
@@
In practice, differences between candidate projection models can be small (Kennedy, 2023), particularly if model classes are nested within one another. In such cases, analysts might have a preference for simpler models, such that more flexible MSMs are selected only if there are clinically meaningful trends in $\chi_{m}(P)$ across calendar time. To formalize this, let $\widehat{\psi}^{*}=\mathop{\mathrm{argmin}}_{\widehat{\psi}_{k}\in\{\widehat{\psi}_{1},\dots,\widehat{\psi}_{K}\}}\widehat{L}(\widehat{\psi}_{k})$ denote the MSM among our set of candidates which minimizes the loss function $\widehat{L}(\widehat{\psi}_{k})$. Let
@@
One can interpret $\epsilon$ as a pooled standard deviation across our best estimate of the calendar time trend, weighting proportional to the size of the eligible population at each trial. Finally, one can choose the simplest model such that $|\widehat{L}(\widehat{\psi}_{k})-\widehat{L}(\widehat{\psi}^{*})|\leq c\epsilon$, for some user-selected constant $c$. There is considerable debate in the literature about how large an effect size needs to be on a standard deviation scale to be clinically meaningful (Tsujimoto et al., 2022), though the notion of clinically meaningful is very application dependent. In simulations, we find that $c=1/4$ works well in our setup, but we present results for a range of other values. Finally, we note that the notion of “simpler model” can be operationalized through any model complexity criterion so long as models can be ordered by this criterion. One reasonable such choice—which we adopt in our application—is ordering candidate MSMs by the dimension of their indexing projection parameter (i.e., the dimension of $\bm{\beta}$).